A right circular cone is inscribed in a cube having a diagonal which
measures 20 square root of 3 cm.

3. What is the surface area of the cone?
4. What is the volume of the space between the cone and the cube?
1
Expert's answer
2020-11-30T12:57:13-0500

The diagonal "d" of the cube with the side "a" is


"d^2=a^2+a^2+a^2=3a^2"

"d=\sqrt{3}a"

Given "d=20\sqrt{3}\ cm."

Then


"a=\dfrac{d}{\sqrt{3}}=\dfrac{20\sqrt{3}}{\sqrt{3}}=20(cm)"

The radius "r" of the circle inscribed in the square with the side "a" is


"r=\dfrac{a}{2}"

Then we have the right circular cone inscribed in a cube with the side "a"


"radius=r=\dfrac{a}{2}, \ height=h=a"

Slant height "L" is


"L=\sqrt{r^2+h^2}=\sqrt{(\dfrac{a}{2})^2+a^2}=\dfrac{\sqrt{5}a}{2}"

3. Surface area of a cone = Base Area + Curved Surface Area of a cone


"A=\pi r^2+\pi rL"

"A=\pi((\dfrac{a}{2})^2+\dfrac{a}{2}(\dfrac{\sqrt{5}a}{2}))=\dfrac{\pi a^2}{4}(1+\sqrt{5})"

"A=\dfrac{\pi (20)^2}{4}(1+\sqrt{5})=100\pi(1+\sqrt{5}) (cm^2)"

"\approx1016.64(cm^2)"

The surface area of the cone is "100\pi(1+\sqrt{5})\ cm^2\approx1016.64\ cm^2."


4. The volume of the cube is


"V_{cube}=a^3"

The volume of the cone is


"V_{cone}=\dfrac{1}{3}\pi r^2h=\dfrac{1}{3}\pi (\dfrac{a}{2})^2 (a)=\dfrac{\pi a^3}{12}"

The volume of the space between the cone and the cube is


"V_{space}=V_{cube}-V_{cone}"

"=a^3-\dfrac{\pi a^3}{12}=\dfrac{ a^3}{12}(12-\pi)"

"V_{space}=\dfrac{(20)^3}{12}(12-\pi)=\dfrac{2000(12-\pi)}{3}(cm^3)"

"\approx5905.605(cm^3)"

The volume of the space between the cone and the cube is

"\dfrac{2000(12-\pi)}{3}\ cm^3\approx 5905.605 \ cm^3."